A Review | Constants An L – R-C series circuit has voltage amplitudes Vư = 180 V, Vc = 120 V, and VR = 160 V. At time t the iPart B At this instant, what is the voltage across the resistor? Express your answer with the appropriate units. .: Å N O 2 ?
Answer #1

The voltage across the resistor (R), inductor (L), and capacitor (L) in series R-L-C circuit is given by

                                       VR= IR sin wt

                                       Vi = IX sin (wt +90°)             ; Xc = iw

                                       VC = IXc sin (at – 90°)             ; Xc = iwc

where \omega is frequency, and t is time.

Given,

                                      Vimar = IX = 180 volt

                                      Vomar = IXc = 120 volt

                                     VRmar = IR = 160 volt

at a certain time t, the voltage across the inductor is 80 volt.

Therefore, we can write

                                       80 = 180 sin(wt +90°)

                               or,   coswt = 80/180

                               or,   coswt = 4/9

             and,                   sin wt = V1 -coswt = V1 -(4/9)2 =

So at that moment of time

Part A:

the voltage across the capacitor is

                                  Vc = 120 sin (wt - 90°) = 120 coswt = 120 X 4/9 = 53.33 volts     

Part B:

the voltage across the capacitor is

                                  VR= 160 sin wt = 160 X V65/9 = 143.33 volts